OpenStudy (anonymous):

fifteen less than 8 times a number is equal to 11 more than 6 times a number.

OpenStudy (anonymous):

8x-15 = 6x+11 8x-6x = 15+11 2x = 26 x = 13

OpenStudy (anonymous):

thanks I don't understand these what so ever